# One-Shot Real-World Demonstration Synthesis for Scalable Bimanual Manipulation

### v1

@@ D.5 VFM-based Object Anchoring in Clutter
Figure 17: Top row: Example cluttered tabletop scenes used to evaluate the perception front-end of BiDemoSyn. Each scene includes the target objects for the pouring task—two bottles (text prompts are blue bottle and green bottle) and two mugs (text prompts blue cup and green cup)—surrounded by visually distracting household items. Two additional tool-like distractors (brush, dustpan) are intentionally introduced to support future extensions such as dual-arm sweeping tasks. Bottom row: High-quality 2D object masks produced by Florence-2 + SAM2 for all queried categories. Despite clutter, color similarity, and semantic ambiguity, the VFM stack consistently isolates each target object with precise boundaries, providing reliable anchors for subsequent pose inference and variant-block adaptation in BiDemoSyn.
Appendix E Summary, Reflection and Outlook
This work presents BiDemoSyn, a framework that bridges the reality gap in bimanual imitation learning by synthesizing diverse, physically feasible demonstrations from a single human example. Through systematic decomposition, vision-guided alignment, and hierarchical optimization, our method eliminates simulation dependencies while scaling data collection by orders of magnitude. Experiments across six contact-rich tasks validate its efficiency (5 seconds per demo), scalability, and generalization to unseen spatial and geometric variations.
However, limitations persist: BiDemoSyn currently assumes quasi-static motions, struggles with highly dynamic tasks (e.g., catching), and faces challenges in extreme shape variations (e.g., deformable objects). Failures primarily stem from orientation estimation errors and grasp-time collisions, highlighting needs for category-agnostic pose estimators and contact-aware motion planners. Future work will extend BiDemoSyn to deformable object manipulation, integrate dynamic perception for moving targets, and explore hybrid learning paradigms (e.g., combining synthetic demos with reinforcement learning). By democratizing real-world demonstration synthesis, this work aims to accelerate progress toward generalizable robotic manipulation.
Appendix F Use of Large Language Models
In preparing this manuscript, we employed the GPT-4o language model strictly for text editing purposes, such as improving readability and polishing grammar. The model was not used to generate research ideas, design methodologies, conduct experiments, or analyze results. All substantive contributions remain entirely those of the authors.

### v2

@@ X-C Cross-Embodiment Transferability Testing
Figure 20: Top: Example cluttered tabletop scenes used to evaluate the perception front-end of BiDemoSyn. Each scene includes the target objects for the pouring task—two bottles (text prompts are blue bottle and green bottle) and two mugs (text prompts blue cup and green cup)—surrounded by visually distracting household items. Two additional tool-like distractors (brush, dustpan) are intentionally introduced to support future extensions such as dual-arm sweeping tasks. Bottom: High-quality 2D object masks produced by Florence-2 + SAM2 for all queried categories.
X-D Statistics and Analysis of Failed Cases
(1) Visual Perception Errors (15%): Inaccurate object segmentation (e.g., mistaking the tabletop as part of the marker pen) leads to misaligned grasps.
(2) Orientation Estimation Errors (32%): Incorrect object pose estimation (e.g., misjudging a spoon’s concave direction) causes malformed trajectories (e.g., flipping to the wrong side).
(3) Alignment Failures (18%): Vision-guided initial frame misalignment (e.g., bottle nozzle offset by >>1cm) propagates errors to subsequent steps (e.g., spilling during pouring).
(4) Initial Grasp Failures (28%): Collisions during pre-grasp motions (e.g., gripper nudging objects) or unstable grasps (e.g., slippage in reorient) prevent task initiation.
(5) Error Propagation (7%): Cumulative errors from earlier stages compound into irreversible failures (e.g., slightly misaligned pluging between pen body and pen cap in plugpen).
X-E VFM-based Object Anchoring in Clutter
To evaluate the applicability of BiDemoSyn under more realistic and visually complex environments, we additionally examine its object-anchoring capability when target objects are placed in cluttered scenes with distractors and partial occlusions. Although the main experiments intentionally isolate environmental variables to study the core challenge of real-world demonstration synthesis, BiDemoSyn is fully compatible with more complex perception conditions due to its reliance on VFMs for object segmentation.
We utilize Florence-2 [89] and SAM2 [74] to detect and segment task-relevant items using either text queries or exemplar-based prompts. Fig. 20 illustrates representative examples where both bottle and mug are correctly segmented out of surrounding tools, utensils, packaging, and nonrigid paper balls. In cluttered tabletop scenes, the VFM consistently produces high-quality masks, even when distractors share similar color similarity and semantic ambiguity. Qualitatively, the predicted masks remain stable across moderate occlusions, enabling the system to isolate the target object with minimal ambiguity and providing reliable anchors for subsequent pose inference and variant-block adaptation in BiDemoSyn.
Because BiDemoSyn’s variant-block adaptation depends only on reliable 2D masks along with centroid- and principal-axis–based orientation inference, accurate object anchoring is the only requirement for extending the system to cluttered environments. Our experiments show that the VFM-based perception pipeline provides robust segmentation to support pose-aligned trajectory modulation and collision-aware grasp adjustments, even when the background is visually nonuniform. However, while VFMs significantly enhance robustness to visual complexity, extreme occlusions or heavy inter-object entanglement may still degrade performance. Extending BiDemoSyn to handle these cases will require multi-view sensing, active view planning, or integrating VLM-driven object-query disambiguation. Nevertheless, these initial results demonstrate that VFM-based anchoring provides a practical and reliable foundation for deploying BiDemoSyn beyond clean tabletop settings, thereby enabling future extensions to household, warehouse, and mobile manipulation scenarios.
XI Summary, Reflection and Outlook
This work introduces BiDemoSyn, a real-world demonstration synthesis framework that addresses a central bottleneck in bimanual imitation learning: how to scale high-quality, physically grounded data collection without reliance on simulation or extensive teleoperation. By decomposing a single kinesthetic demonstration into invariant coordination structures and object-dependent adaptations, and by coupling vision-guided alignment with lightweight trajectory optimization, BiDemoSyn enables the rapid generation of diverse, feasible bimanual demonstrations directly in the physical world. Extensive experiments across six contact-rich tasks demonstrate its efficiency, scalability, and strong generalization to novel object poses and geometries.
Beyond the original one-shot setting, we further show that BiDemoSyn naturally extends to few-shot demonstration synthesis, where additional demonstrations primarily enhance out-of-distribution generalization without altering the core pipeline. Moreover, we demonstrate zero-shot cross-embodiment policy deployment, where policies trained on BiDemoSyn data transfer effectively to a distinct humanoid-style dual-arm robot, preserving high success rates across representative tasks. These results highlight the importance of object-centric observations and embodiment-agnostic action representations, and suggest that BiDemoSyn supports not only scalable data generation, but also robust policy reuse across hardware platforms.
Despite these advances, several limitations remain. BiDemoSyn currently assumes quasi-static task execution and rigid or semi-rigid objects, making highly dynamic interactions and deformable object manipulation challenging. Failure cases are primarily attributed to perception noise—particularly orientation estimation for small or symmetric objects—and occasional contact inaccuracies during grasping and insertion. Addressing these issues will likely require tighter perception–control coupling, category-agnostic orientation reasoning, and more expressive contact-aware planners.
Looking forward, we see multiple promising directions. Future work will explore extending BiDemoSyn to deformable and articulated objects, incorporating dynamic and multi-view perception, and integrating closed-loop feedback into the synthesis pipeline. More broadly, we envision BiDemoSyn as a foundation for hybrid data-centric learning paradigms, combining one-shot, few-shot, and synthesized demonstrations to balance data efficiency with maximal generalization. By lowering the barrier to acquiring large-scale, real-world bimanual data, this work aims to accelerate progress toward practical, general-purpose robotic manipulation.